# Perturbation-Robust Predictive Modeling of Social Effects by Network Subspace Generalized Linear Models

### v4

@@ 4.2 Network perturbations by from deep-learning-based embedding methods
We now consider another application scenario in which the proposed model can be used. Suppose we want to use embedding methods from deep-learning communities to extract the network information. Multiple recent works (Pozek et al., 2019; Pranathi and Prathibhamol, 2021; Liu and Huang, 2024) take this strategy to incorporate network information, with the belief that these methods can capture high-order network relations more effectively by their highly nonlinear operations.
Our subspace generalized linear model, with its flexibility in specifying a proper subspace $S_{K}(P)$, can seamlessly leverage this embedding information. Specifically, we can assume the Euclidean similarities of the embedded vectors as the perturbed relational information $\hat{P}$, with the true relational matrix being an unobserved similarity matrix that can be different from the random embedded similarities. In these cases, even if the network is usually treated as fixed, the embedding algorithms are typically random by nature. This randomness in embeddings raises concerns about the validity of modeling and inference if one uses a specific embedding in the model. In this section, we use simulation experiments to evaluate the validity of our model’s inference under such permutations of embeddings. The study of statistical properties of the embedding methods is rare in literature. The only one we know is from Zhang and Tang (2023) on community detection. To our knowledge, the current experiment is the first empirical study in the literature to investigate the deep-learning-based network embedding methods’ impacts on downstream statistical inference tasks.
We consider three popular network embedding methods, DeepWalk (Perozzi et al., 2014), Node2Vec (Grover and Leskovec, 2016), and Diff2Vec (Rozemberczki and Sarkar, 2018) to demonstrate these scenarios. DeepWalk was one of the earliest graph embedding methods from the deep learning community, and Node2Vec is a generalization of DeepWalk. Diff2Vec uses the more recent diffusion framework to define the embeddings. The implementations of DeepWalk and Node2Vec are available in the Python package node2vec (Grover and Leskovec, 2016), and Diff2Vec is implemented in the Python package karateclub (Rozemberczki et al., 2020). In our simulation, we always use the recommended configurations of these methods. For DeepWalk and Node2Vec, each embedding is based on $10$ walks per node of length $80$. For Node2Vec, the return probability is set to $0.5$. For Diff2Vec, we use $20$ trees per node of size $80$. The network embedding dimension is always set to $3$.
Design of relational matrix. We first generate a network $A$ from one of the three models in Section 4.1, and fix the network $A$. Given $A$, all three embedding methods are random and result in different embeddings each time. Therefore, for each embedding method, suppose $\mathcal{F}$ is the embedding of $A$ and, intuitively, we can use $\mathcal{F}\mathcal{F}^{\top}$ as the available similarity matrix we can use from data. The perturbation of network information comes from the randomness of $\mathcal{F}$. Specifically, in this context, we set the true relational matrix as the oracle central similarity $P=\mathbb{E}[\mathcal{F}\mathcal{F}^{\top}]$, and $\hat{P}=\mathcal{F}\mathcal{F}^{\top}$. The design matrix $X$ and other quantities are generated in the same manner as in Section 4.1 based on the current $P$.
@@
Table 5: Median MSE, coverage probability for subspace logistic regression with different types of network of size $2000$ under network embedding perturbations.
@@
Table 6: Median MSE, coverage probability for subspace Poisson regression with different types of network of size $2000$ under network embedding perturbations.
@@
Table 5 and Table 6 illustrate all the performance metrics under the perturbation of different embedding algorithms on the three network types with varying average degrees when $n=2000$. Note that the previous benchmark method, RNC, is not applicable in this new setting and was removed from the comparison. This also demonstrates the flexibility of our subspace-based model.
For each of the three embedding mechanisms, the estimation accuracy and inference correctness (measured by the coverage probability) exhibit mild improvements with the increase of network density. Among the three mechanisms, Diff2Vec is more sensitive to network density. For example, on networks with an average degree of $2\log n$, the coverage probability misses the target level by about 5%. DeepWalk and Node2Vec are more robust. Overall, the impact of network density is weaker on all embedding mechanisms compared to the random network perturbation mechanisms. The most important observation from these experimental results is that our model estimation remains accurate, and the statistical inference is still approximately correct unless the network is very sparse. This result demonstrates the applicability of our inference framework in broader scenarios in modern machine learning.
@@ 5 Social and Educational Effect Study of School Conflicts
With the proposed model, we will analyze data from the school conflict study introduced in Section 1.1. Following the original strategy of Paluck et al. (2016), we use the allocation of orange wristbands to examine the impact of anti-conflict interventions on behaviors that promote a positive school climate, rewarding actions that are friendly or conflict-mitigating. If a student receives an orange wristband, the response variable $Y$ is set to $1$; otherwise, it is 00. We employ a logistic model to analyze this response.
To identify features strongly associated with the allocation of orange wristbands, we use individual attributes from the supplementary materials of Paluck et al. (2016) as potential predictors. These include Treatment (participation in weekly training: Yes/No), Gender (Male/Female), Race (White/Hispanic/Black/Asian/Others), Grade, “Friends-like-house” (friends say I have a nice house: Yes/No), and Home-language (speaks another language at home: Yes/No). Additionally, we incorporate GPA (grade point average on a 4.0 scale) and a binary covariate, Influencer (nominated by the teacher as influential). A school-specific effect parameter is included for each school to account for school-level differences. We use the largest connected components from each school, removing students with incomplete information. After data processing, we have data on 8685 students from 25 schools, with 1391 students receiving an orange wristband. Based on our evaluation, embedding methods such as Node2Vec do not improve predictive performance (see Section A.3) for the current dataset. Therefore, we will directly use the observed networks for better interpretability and to align with the original analysis of Paluck et al. (2016).
@@ 5.1 Model fitting and interpretations
We use the average of two friendship adjacency matrices from two survey waves (at the start and end of the school year) as our weighted adjacency matrix to measure the relations between students. It turns out that using either matrix alone yields the same analysis, thanks to the robustness of our framework to network perturbations (see Section A.2). This is a crucial advantage of our framework.
@@
Under our model, using the full dataset, the $\chi^{2}$ test for network effects gives a very small p-value ($<10^{-8}$), indicating a significant contribution of the network information. The estimated $\hat{\alpha}_{i}$’s are shown in Figure 2. The corresponding estimated coefficients of covariates are included in Section A in Appendix. The network effects show considerable variation in magnitude across different schools, with strong social impacts observed in a few schools, while in most others, the social network impact is minor. Consequently, we focus our analysis on the schools with strong network effects for deeper insights.
We define a school-specific network-effect-strength $t_{j}:=\sum_{i\in O_{j}}|\hat{\alpha}_{i}|/\sum_{i\in O_{j}}|x_{i}^{\top}\hat{\beta}|$, where $O_{j}$ is the index of students in the $j$th school. We select the five schools with the largest $t_{j}$ values (Schools 1, 22, 27, 31, and 48) for further analysis. We then apply our SP logistic model, standard logistic regression, and the RNC logistic regression to the data. For our model and the standard logistic regression, important predictors are selected using backward elimination, whereby variables with the largest p-values exceeding 0.05 after Bonferroni correction are removed sequentially until no further elimination is possible. As the RNC model lacks an inference framework, we retain all variables in that model. The results of the three fitted models and their corresponding p-values are summarized in Table 7.222Note that the p-values do not account for the selection of the five schools based on data and the backward elimination of variables. These results are used primarily for qualitative interpretations. In Section 5.2, we validate the models more rigorously through prediction performance, with the randomness of the selection procedure accounted for. School-specific effects are not included in the table.
Again, our $\chi^{2}$ test gives a very small p-value, providing strong evidence of social effects. The coefficient for Treatment is estimated similarly by all three models. This is expected since, in the study, whether a student is enrolled in the training is randomly determined by the experiment design, making this variable uncorrelated with any other effects. However, unlike the RNC, our method and the standard logistic regression can use their p-values to show that the treatment effect is indeed significant. This finding aligns with the analysis of Paluck et al. (2016), which used a more sophisticated causal inference framework to study the data with the design information.
Table 7: Model fitting and inference results on five schools with strongest effect schools.
@@
Treatment 0.644 $<10^{-3}$ 0.629 $<10^{-3}$ 0.684
Gender:Male -0.450 $<10^{-3}$0.568 $<10^{-3}$0.521
@@
Network Effect 8.5×$\times 10^{-3}$
Since the RNC does not provide inference and variable selection, we focus on comparing our method with standard logistic regression. The two models yield very different inferences for the effects of Gender, Grade, and Race. Notably, Gender is the only predictor besides Treatment that remains in the final selection based on our model. The standard logistic regression estimates a 25% stronger gender effect and finds statistically significant negative effects for Grade and Race. The main difference between our model and standard logistic regression is the inclusion of network effects, suggesting that the differential predictors may be cohesive according to network structures. This phenomenon is intuitively reasonable. For example, students are more likely to be friends with others in the same grade. We can empirically verify these conjectures. Figure 3 shows the gender, grade, and race information in one of the five schools: students tend to befriend others of the same gender, grade, and race. Similar patterns can be observed in other schools (see Figure 7 in Appendix A). Therefore, these predictors exhibit network cohesion, explaining the differential results between our method and standard logistic regression.
@@
In summary, we have shown that social network information give important impacts in the current problem. Though both the RNC and our SP model can incorporate network information in building the logistic regression model, the available inference framework in our model provides a substantial advantage in understanding the data with more conclusive insights: both the social effect and the conflict-mitigating training are statistically significant in this example. Compared with the standard logistic regression, all the qualitative differences in estimated effects can be explained by the network cohesion phenomenon, which can be empirically verified.
All previous discussions focus on model interpretation and we have seen that differences between our model and the standard logistic regression are reasonable. Next, we use prediction performance to validate the effectiveness of our model compared with the standard logistic regression.

### v6

@@ 4.2 Network perturbations from deep-learning-based embedding methods
We now consider another application scenario in which the proposed model can be used. Suppose we want to use embedding methods from deep-learning community to extract the network information. Multiple recent works (Pozek et al., 2019; Pranathi and Prathibhamol, 2021; Liu and Huang, 2024) take this strategy to incorporate network information, with the belief that these methods can capture high-order network relations more effectively by their highly nonlinear operations.
@@
Table 7: Median MSE ($\times 10^{2}$), coverage probability and MSPE ($\times 10^{2}$) for subspace logistic regression with different types of network of size $1000$ under network embedding perturbations.
Method Network avg. degree MSE Coverage MSPE
Our Model Logistic Reg
DeepWalk $2\log n$ 0.60 94.6% 0.12 0.61
SBM $\sqrt{n}$ 0.59 94.9% 0.11 1.75
$n^{2/3}$ 0.58 94.9% 0.11 2.04
$2\log n$ 0.75 94.8% 0.13 0.34
DCBM $\sqrt{n}$ 0.62 94.5% 0.13 1.98
$n^{2/3}$ 0.60 94.8% 0.12 1.31
$2\log n$ 0.77 95.2% 0.08 1.08
Diag $\sqrt{n}$ 0.58 95.1% 0.09 1.56
$n^{2/3}$ 0.62 95.2% 0.09 1.54
Node2Vec $2\log n$ 0.59 94.6% 0.12 1.61
SBM $\sqrt{n}$ 0.59 94.9% 0.12 1.35
$n^{2/3}$ 0.60 94.9% 0.12 1.90
$2\log n$ 0.69 94.8% 0.12 0.29
DCBM $\sqrt{n}$ 0.68 94.9% 0.12 0.24
$n^{2/3}$ 0.65 94.8% 0.12 1.11
$2\log n$ 0.56 94.9% 0.09 1.45
Diag $\sqrt{n}$ 0.52 95.0% 0.08 1.25
$n^{2/3}$ 0.58 94.9% 0.09 1.31
Diff2Vec $2\log n$ 0.56 94.9% 0.09 1.12
SBM $\sqrt{n}$ 0.52 95.0% 0.09 1.08
$n^{2/3}$ 0.58 94.9% 0.09 1.26
$2\log n$ 0.66 94.3% 0.14 1.18
DCBM $\sqrt{n}$ 0.72 94.7% 0.12 1.21
$n^{2/3}$ 0.62 94.8% 0.11 1.19
$2\log n$ 0.56 94.9% 0.09 0.86
Diag $\sqrt{n}$ 0.52 95.0% 0.09 1.01
$n^{2/3}$ 0.58 94.9% 0.09 1.09
Our subspace generalized linear model, with its flexibility in specifying a proper subspace $S_{K}(P)$, can seamlessly leverage this embedding information. Specifically, we can assume the inner product similarities of the embedded vectors as the perturbed relational information $\hat{P}$, with the true relational matrix being an unobserved similarity matrix that can be different from the random embedded similarities. In these cases, even if the network is usually treated as fixed, the embedding algorithms are typically random by nature. This randomness in embeddings raises concerns about the validity of modeling and inference if one uses a specific embedding in the model. In this section, we use simulation experiments to evaluate the validity of our model’s inference under such perturbations of embeddings. The study of statistical properties of the embedding methods is rare in the literature. To our knowledge, Zhang and Tang (2023) provides related analysis for community detection; we are not aware of prior empirical studies examining how deep‑learning–based embeddings affect downstream inference.
We consider three popular network embedding methods, DeepWalk (Perozzi et al., 2014), Node2Vec (Grover and Leskovec, 2016), and Diff2Vec (Rozemberczki and Sarkar, 2018) to demonstrate these scenarios. DeepWalk was one of the earliest graph embedding methods from the deep learning community, and Node2Vec is a generalization of DeepWalk. Diff2Vec uses the more recent diffusion framework to define the embeddings. The implementations of DeepWalk and Node2Vec are available in the Python package node2vec (Grover and Leskovec, 2016), and Diff2Vec is implemented in the Python package karateclub (Rozemberczki et al., 2020). In our simulation, we always use the recommended configurations of these methods. For DeepWalk and Node2Vec, each embedding is based on $10$ walks per node of length $80$. For Node2Vec, the return probability is set to $0.5$. For Diff2Vec, we use $20$ trees per node of size $80$. The network embedding dimension is always set to 33.
Design of relational matrix. We first generate a network $A$ from one of the three models in Section 4.1, and fix the network $A$. Given $A$, all three embedding methods are random and result in different embeddings each time. Therefore, for each embedding method, suppose $\mathcal{F}$ is the embedding of $A$ and, intuitively, we can use $\mathcal{F}\mathcal{F}^{\top}$ as the available similarity matrix from data. The perturbation of network information comes from the randomness of $\mathcal{F}$. Specifically, in this context, we set the true relational matrix as the oracle central similarity $P=\mathbb{E}[\mathcal{F}\mathcal{F}^{\top}]$, and $\hat{P}$$\mathcal{F}\mathcal{F}^{\top}$. The design matrix $X$ and other quantities are generated in the same manner as in Section 4.1 based on the current $P$.
Table 8: Median MSE ($\times 10^{2}$), coverage probability and MSPE ($\times 10^{2}$) for subspace logistic regression with different types of network of size $2000$ under network embedding perturbations.
@@
Table 9: Median MSE ($\times 10^{2}$), coverage probability, and MSPE for subspace Poisson regression with different types of network of size $1000$ under network embedding perturbations.
Method Network avg. degree MSE Coverage MSPE
Our Method Poisson Reg
DeepWalk $2\log n$ 0.14 93.3% 2.12 12.1
SBM $\sqrt{n}$ 0.12 94.5% 2.02 62.7
$n^{2/3}$ 0.10 94.5% 1.92 56.7
$2\log n$ 0.16 93.5% 2.97 23.0
DCBM $\sqrt{n}$ 0.09 92.9% 3.45 59.9
$n^{2/3}$ 0.12 94.3% 1.83 21.7
$2\log n$ 0.20 94.9% 1.12 11.7
Diag $\sqrt{n}$ 0.11 94.8% 0.88 24.3
$n^{2/3}$ 0.13 94.9% 0.84 25.0
Node2Vec $2\log n$ 0.12 93.0% 3.09 64.7
SBM $\sqrt{n}$ 0.13 94.2% 1.84 19.4
$n^{2/3}$ 0.13 94.3% 1.97 43.5
$2\log n$ 0.09 93.7% 3.74 27.8
DCBM $\sqrt{n}$ 0.09 93.9% 2.37 4.33
$n^{2/3}$ 0.08 93.7% 3.58 42.1
$2\log n$ 0.16 94.7% 0.62 12.7
Diag $\sqrt{n}$ 0.10 95.0% 0.80 13.7
$n^{2/3}$ 0.11 94.7% 0.89 22.2
Diff2Vec $2\log n$ 0.16 90.9% 2.30 25.5
SBM $\sqrt{n}$ 0.15 93.9% 1.72 16.0
$n^{2/3}$ 0.15 94.4% 1.33 27.3
$2\log n$ 0.15 91.9% 3.12 27.7
DCBM $\sqrt{n}$ 0.18 93.5% 2.19 24.5
$n^{2/3}$ 0.11 93.9% 2.14 20.2
$2\log n$ 0.16 90.9% 2.30 6.44
Diag $\sqrt{n}$ 0.27 94.9% 0.92 9.21
$n^{2/3}$ 0.18 94.9% 0.79 12.1
Table 10: Median MSE ($\times 10^{2}$), coverage probability and MSPE for subspace Poisson regression with different types of network of size $2000$ under network embedding perturbations.
@@
Tables 7 to 10 present the performance metrics under perturbations from different embedding algorithms, evaluated across three network types with varying average degrees for sample sizes n=1000, 2000n=1000,\ 2000. Additional results for $n=500$ are provided in Section I. Note that the previous benchmark method, RNC, is not applicable in this new setting and was removed from the comparison. This also demonstrates the flexibility of our subspace-based model.
For each of the three embedding mechanisms, the estimation accuracy and inference correctness (measured by the coverage probability) exhibit mild improvements with the increase of network density. But overall, the density is no longer a very clear indicator of the perturbation level in this case, because the perturbation is contributed by the randomness of the embedding algorithms. Among the three mechanisms, Diff2Vec is more vulnerable to density change, and overall results in larger perturbations. For example, on networks with an average degree of $2\log n$ for sample size $n=2000$, the coverage probability misses the target level by a lot. DeepWalk and Node2Vec are more robust and the resulting perturbations tend to satisfy the small projection perturbation requirements. For embedding methods, our model estimation remains accurate, and the statistical inference is still approximately correct. This result demonstrates the applicability of our inference framework in broader scenarios in modern machine learning.
@@ 5 Social and Educational Effect Study of School Conflicts
With the proposed model, we will analyze data from the school conflict study introduced in Section 1.1. Following the original strategy of Paluck et al. (2016), we use self-reported wearing of an orange wristband to assess the impact of anti-conflict interventions on behaviors that promote a positive school climate: Each week, orange wristbands were distributed to students who were observed engaging in positive, conflict‑reducing behavior. All students in the schools were eligible to receive a wristband as recognition for the conflict-mitigating behaviors. If a student is reported wearing an orange wristband, the response variable $Y$ is set to $1$; otherwise, it is 0. We fit the subspace logistic regression described in Section 2 to analyze this response.
To identify features strongly associated with the allocation of orange wristbands, we use individual attributes from the supplementary materials of Paluck et al. (2016) as potential predictors. These include Treatment (participation in weekly training: Yes/No), Gender (Male/Female), Race (White/Hispanic/Black/Asian/Others), Grade, “Friends-like-house” (friends say I have a nice house: Yes/No), and Home-language (speaks another language at home: Yes/No). Additionally, we incorporate GPA (grade point average on a 4.0 scale) and a binary covariate, Influencer (nominated by the teacher as influential). An individual school effect parameter is introduced for each school to account for school-level differences. We remove students with missing values and then use the largest connected component from each school. The final dataset contains 8,685 students from 25 schools. Each network has an average size of 347.1 students and an average degree of 10.7. Among all students, 1,391 received an orange wristband.
Based on our evaluation, embedding methods mentioned before, such as Node2Vec, do not improve predictive performance (see Section J) for the current dataset. Therefore, we will directly use the observed networks for better interpretability.
@@ 5.1 Model fitting and interpretations
We use the average of two friendship adjacency matrices from two survey waves (at the start and end of the school year) as our denoised $\hat{P}$ to measure the relations between students. It turns out that in this example using either matrix alone yields similar analyses, thanks to the robustness of our framework to network perturbations (see Section A.2). This robustness is a crucial advantage of our framework.
@@
Our model incorporates the top 31 eigenvectors of the adjacency matrix to capture network effects, selected by Chatterjee (2015). The $\chi^{2}$ test for network effects gives a very small $p$-value ($10^{-8}$), indicating a significant contribution of the network information. The estimated $\hat{\alpha}_{i}$’s are shown in Figure 2. The corresponding estimated coefficients of covariates are included in Section A. The estimated value of $r$ is 0, suggesting that there is no overlap between the covariate and the network structure.
Based on Figure 2, network effects vary significantly in magnitude between different schools. A few of the schools exhibit strong social influences, while many other schools exhibit minor social network effects. Aggregating all schools together would dilute the significance of social effects. To gain a more comprehensive understanding of these dynamics, we introduce another layer of analysis on the schools with the most pronounced network effects, allowing us to explore the underlying factors driving these stronger social influences.
[…]
In the selected dataset of five schools, our model picks $K=13$ while $r$ is estimated to be 0. Again, the $\chi^{2}$ test gives a very small $p$-value, providing strong evidence of social effects. To better interpret the estimated network effect $\alpha$, we compute the correlation between $\left|\alpha\right|$ and a binary indicator of seed-eligible students, identified using the algorithm described in the supplement of Paluck et al. (2016), along with four network centrality metrics: degree centrality, betweenness centrality, eigenvector centrality, closeness centrality. The results are presented in Table 13.
[…]
The estimated treatment coefficient is similar across the three models. This might be expected due to the random assignment implemented by the experimenters, making this variable uncorrelated with other effects. However, unlike the RNC, our method and the standard logistic regression can use their $p$-values to show that the treatment effect is indeed significant.
Table 11: Model fitting and inference results (before variable selection through backward elimination) on the five schools with the strongest network effects.
@@
Treatment 0.636 0.005 0.678 0.004 0.684
Gender: Male -0.454 0.002 -0.440 0.003 -0.521
Grade 0.131 0.141 -0.119 0.126 -0.253
Friends like house 0.011 0.472 -0.066 0.650 -0.077
Home language 0.203 0.137 0.271 0.134 0.212
GPA 0.106 0.245 0.054 0.721 -0.230
Influencer 0.393 0.047 0.366 0.092 0.504
Race: White -0.369 0.058 -0.296 0.192 -0.488
Race: Black -0.085 0.401 -0.102 0.756 -0.229
Race: Hispanic -0.457 0.025 -0.512 0.024 -0.544
Race: Asian 0.250 0.217 0.260 0.404 0.158
Network Effect $6.0\times 10^{-3}$
School 22 -1.447 0.002 -1.964 $10^{-3}$
School 27 0.960 0.018 -0.002 0.993
School 31 -0.033 0.470 -0.010 0.965
School 48 -0.455 0.164 0.065 0.784
Table 12: Model fitting and inference results (after variable selection through backward elimination) on the five schools with strongest network effects.
Our Model Logistic Reg RNC
coef. $p$-value coef. $p$-value coef. $p$-value
Treatment 0.644 $10^{-3}$ 0.629 $10^{-3}$ 0.684
Gender: Male -0.450 $10^{-3} -$0.568 $10^{-3} -$0.521
@@
[…]
School 22 -1.281 $10^{-3} -$2.101 $10^{-3}$
School 27 1.003 0.002
Table 13: Correlation between $\left|\alpha\right|$ and degree, betweenness, eigenvector, and closeness centrality, and seed eligibility, for the full dataset and for schools with strong network effects.
Degree Cen Betweenness Cen Eigenvector Cen Closeness Cen Seed Eligibility
Full dataset 0.258 0.112 -0.043 -0.025 -0.003
Selected dataset 0.300 0.199 0.477 -0.189 -0.031
Since the RNC does not provide inference or variable selection, we focus on comparing our method with standard logistic regression. The two models yield very different inferences for the effects of Gender, Grade, and Race. Notably, Gender is the only predictor besides Treatment that remains in the final selection based on our model. The standard logistic regression estimates a 25% stronger gender effect and finds statistically significant negative effects for Grade and Race. The main difference between our model and standard logistic regression is the inclusion of network effects, suggesting that the differential predictors may be cohesive according to network structures. This phenomenon is intuitively reasonable. For example, students are more likely to be friends with others in the same grade. We can empirically verify these conjectures. Figure 3 shows the gender, grade, and race information in one of the five schools: students tend to befriend others of the same gender, grade, and race. Similar patterns can be observed in other schools (see Figure 7 in Section Al). Therefore, these predictors exhibit network cohesion, explaining the differential results between our method and standard logistic regression.
To further support the statement “students tend to befriend others of the same gender, grade, and race” quantitatively, we include the following summary table. It reports, for each attribute, the proportion of same-attribute friendships, the expected proportion under random mixing, and the assortativity coefficient. The assortativity coefficient proposed in Newman (2003) is defined as the Pearson correlation between the attribute values:
$r=\frac{\sum_{i}e_{ii}-\sum_{i}a_{i}b_{i}}{1-\sum_{i}a_{i}b_{i}},$
where $e_{ij}$ denotes the fraction of edges connecting a node of type $i$ to a node of type $j$, and $a_{i}=\sum_{j}e_{ij}$, $b_{j}=\sum_{i}e_{ij}$ represent the fraction of edges attached to nodes of type $i$ and $j$, respectively. In the case of undirected networks, $e_{ij}=e_{ji}$ and $a_{i}=b_{i}$.
An assortativity of $r=1$ corresponds to perfect homophily, $r=0$ to random mixing, and $r<0$ to disassortative mixing. We estimated its standard error using a jackknife procedure, by sequentially removing each edge, recalculating the assortativity $r_{i}$, and computing the variance as
$s^{2}_{r}:=\sum_{i=1}^{M}\left(r_{i}-r\right)^{2},$
where $M$ is the number of edges and $r$ is the observed assortativity. The standard error $s_{r}$ is the square root of this sum. The results are summarized below:
Table 14: Observed and expected proportions of same-attribute friendships and assortativity coefficients based on Gender, Grade, and Race in School 1’s Friendship Network
Same-Attribute Edges Proportion Assortativity Statistics
Observed Expected Assortativity Std. Error
Gender 76.3% 50.1% 0.524 0.021
Grade 88.5% 33.5% 0.827 0.012
Race 49.0% 42.3% 0.122 0.019
All three covariates exhibit significant assortative mixing, consistent with the discussion in Newman (2003), with the effect being especially strong for gender and grade. This confirms that these covariates are highly correlated with the network structure, which helps explain the different coefficient magnitudes observed between our model and the standard logistic regression. Although race has the lowest assortativity among the three, the effect is still highly statistically significant (with a value more than six standard deviations from zero). This finding aligns with our observation that our model and the standard logistic regression differed in their variable selection for race only for the selected schools, but not in the full dataset.
@@
The positive effect of the workshop observed in our analysis is consistent with the main conclusion of the original study by Paluck et al. (2016). However, we emphasize that this agreement is only at a high level. Our analysis differs from that of Paluck et al. (2016) in several important respects. First, Paluck et al. (2016) focus on a subpopulation of students who are connected to at least one potentially treated peer, whereas our analysis does not impose such a restriction and includes all available observations from treated schools. Second, their analysis is explicitly causal in nature, relying on the original randomized design and causal inference methods. In contrast, our results are descriptive and inferential but do not carry a causal interpretation.
In summary, we have shown that social network information has important impacts in the current problem. Though both the RNC and our model can incorporate network information in building the logistic regression model, the available inference framework in our model provides a substantial advantage in understanding the data with more conclusive insights: both the social effect and the conflict-mitigating training are statistically significant in this example. Compared with the standard logistic regression, all the qualitative differences in estimated effects can be explained by the network cohesion phenomenon, which can be empirically verified.
All previous discussions focus on model interpretation and we have seen that differences between our model and the standard logistic regression are reasonable. Next, we use prediction performance to validate the effectiveness of our model compared to the standard logistic regression.